# Stochastic heat flow by moments

### v1

@@ 3.1. Lindeberg’s method
We use Lindeberg’s exchange method. Divide $[0,1]$ into $N$ intervals of equal length, with $t_{i}:=i/N$, fix any $L^{2}$-applicable mollifier $\{u_{\ell}\}_{\ell}$, and let
$h_{i}(\ell,N)$ $:=r_{i-1}(\ell,N)-r_{i}(\ell,N)\ ,\qquad r_{i}(\ell,N):=\mathbf{E}\big{[}h\big{(}\widetilde{Z}_{0,t_{i}}\bullet_{u_{\ell}}Z_{t_{i},1}f\big{)}\big{]}\ .$ (3.2)
We take $Z$ and $\widetilde{Z}$ to be independent within expectations like the one above. The product $\bullet_{u_{\ell}}$ above is well-defined thanks to Lemma 2.3. We then write
$\mathbf{E}\big{[}h(Z_{0,1}f)\big{]}-\mathbf{E}\big{[}h(\widetilde{Z}_{0,1}f)\big{]}=h_{1}(\ell,N)+\ldots+h_{N}(\ell,N)\ ,$ (3.3)
and will send $\ell\to\infty$ first and $N\to\infty$ later.
We seek to derive an approximation of $r_{i-1}(\ell,N)$ via the Taylor expansion. First, by Corollary 2.5(b) for $T=\{t_{i-1},t_{i},1\}$, $r_{i-1}(\ell,N)$ is equal to $\lim_{\ell^{\prime}\to\infty}\mathbf{E}[h\big{(}\widetilde{Z}_{0,t_{i-1}}\bullet_{u_{\ell}}Z_{t_{i-1},t_{i}}\bullet_{u_{\ell^{\prime}}}Z_{t_{i},1}f)]$. Further using Proposition 2.4 with $T=\{0,t_{i-1},t_{i},1\}$ and $S=\{t_{i-1},t_{i}\}$ gives
$\lim_{\ell\to\infty}r_{i-1}(\ell,N)=\lim_{(\ell,\ell^{\prime})\to(\infty,\infty)}\mathbf{E}\big{[}h\big{(}\widetilde{Z}_{0,s}\bullet_{u_{\ell}}Z_{s,t}\bullet_{u_{\ell^{\prime}}}Z_{t,1}f)\big{]}\Big{|}_{(s,t)=(t_{i-1},t_{i})}\ .$ (3.4)
@@
$\limsup_{\ell\to\infty}\Big{|}r_{i-1}(\ell,N)-\sum_{n=0}^{3}\frac{1}{n!}h^{n}_{t_{i-1},t_{i}}\Big{|}\leq c(h)\,v_{t_{i-1},t_{i}}\ .$ (3.8)
Next, we claim that (3.8) holds also for $r_{i}(\ell,N)$ in place of $r_{i-1}(\ell,N)$. Indeed, running the same argument above shows that (3.4)–(3.6) hold for $r_{i}(\ell,N)$ in place of $r_{i-1}(\ell,N)$, for $W_{s,t}$ in place of $\widetilde{W}_{s,t}$, and with $\ell$ and $\ell^{\prime}$ swapped. The swapping of $\ell$ and $\ell^{\prime}$ does not matter because the second limit in (3.4) is taken jointly in $\ell,\ell^{\prime}$. When evaluating the expectations in (3.6)–(3.7), we used only the moment formula (2.18) and the independence of $\widetilde{Z}_{0,s}$, $W_{s,t}$, $Z_{t,1}$. These properties hold also for $\widetilde{W}_{s,t}$ in place of $W_{s,t}$. Hence (3.8) holds for $r_{i}(\ell,N)$ in place of $r_{i-1}(\ell,N)$.
Let us bound (3.3). Combining (3.8) for $r_{i-1}(\ell,N)$ and for $r_{i}(\ell,N)$ gives
$\limsup_{\ell\to\infty}|h_{i}(\ell,N)|\leq c(h)\,v_{t_{i-1},t_{i}}\ .$ (3.9)
Next, in (3.7), use the symmetry of $\mathcal{Q}^{[\![4]\!]}(s)$ (its generator is self-adjoint) to write the inner product as $\langle\mathcal{Q}^{[\![4]\!]}(s)\bigotimes_{i=1}^{4}g_{i},\mathcal{W}^{[\![4]\!]}(t-s)\mathcal{Q}^{[\![4]\!]}(1-t)\bigotimes_{i=1}^{4}g^{\prime}_{i}\rangle$, and, for $p\in(2,\infty)$, use (2.16) to bound $\|\mathcal{Q}^{[\![4]\!]}(s)\bigotimes_{i=1}^{4}g_{i}\|_{p,1}$ and $\|\mathcal{Q}^{[\![4]\!]}(1-t)\bigotimes_{i=1}^{4}g^{\prime}_{i}\|_{p,1}$ by $c=c(p,\Gamma)$, where we encoded the dependence on $g_{i},g^{\prime}_{i}$ by $\Gamma$. The result gives $v_{s,t}\leq c(p,h,\Gamma)\|\mathcal{W}^{[\![4]\!]}(t-s)\|_{p,1\to q,1}$, where $11/p+1/q=$ 1. Inserting this bound into (3.9) and summing the result over $i=1,\ldots,N$ give, for every $p\in(2,\infty)$,
$\big{|}\mathbf{E}\big{[}h(Z_{0,1}f)\big{]}-\mathbf{E}\big{[}h(\widetilde{Z}_{0,1}f)\big{]}\big{|}\leq c(p,h,\Gamma)\,N\big{\|}\mathcal{W}^{[\![4]\!]}(\tfrac{i}{N})\big{\|}_{p,1\to q,1}\ .$ (3.10)
The proof is completed upon applying Lemma 3.1 with $42n=$ 4 and any $p\in(6,\infty)$ and sending $N\to\infty$. We prove the lemma in the next section.
@@
For any $2n\in 2\mathbb{Z}_{>0}$, $p\in(2,\infty)$, and $a\in(0,\infty)$, with $11/p+1/q=$ 1,
@@
The key to the proof is to derive a useful form of $\mathcal{W}^{[\![2n]\!]}(t)$. Fix $2n\in 2\mathbb{Z}_{>0}$, $p\in(2,\infty)$, and $a\in(0,\infty)$, write $c=c(n,p,a)$, and take $11/p+1/q=$ 1. Recall the expression of $\mathcal{W}^{[\![2n]\!]}(t)$ from (2.20) and also (2.12). The key to the proof is to rewrite that expression in a new form that exposes the nature of $\mathrm{Dgm}_{*}[\![2n]\!]$. To state the new form, we need some notation. For $\alpha\in\mathrm{Pair}(\omega)$, let
@@
$\mathcal{B}^{\ell}=\mathcal{B}^{\ell}_{\vec{\eta}}(\vec{s})=\mathcal{B}^{\ell}_{\eta_{1},\ldots,\eta_{\ell}}(s_{1},\ldots,s_{\ell}):L^{2}(\mathbb{R}^{2}\times\mathbb{R}^{2\omega_{\ell}\setminus\eta_{\ell}})\to L^{q,a}(\mathbb{R}^{2\omega_{\ell}})$ (3.14)
@@
Finally, define $\mathcal{B}_{\alpha}(t):L^{2}(\mathbb{R}^{2[\![2n]\!]})\to L^{q,a}(\mathbb{R}^{2[\![2n]\!]})$,
$\mathcal{B}_{\alpha}(t):=\sum_{\vec{\eta}\in\mathrm{Dgm}_{\uparrow}[\![2n]\!],\eta_{n}=\alpha}\int_{\Sigma(t)}\mathrm{d}\vec{s}\,\mathcal{B}^{n}_{\vec{\eta}}(s_{1},\ldots,s_{n})\otimes\mathcal{P}^{[\![2n]\!]\setminus(\eta_{1}\cup\cdots\cup\eta_{n})}(s_{1}+\ldots+s_{n})\ .$ (3.18)
We will prove in the next few paragraphs that $\mathcal{W}^{[\![2n]\!]}(t)$ in (2.20) is equal to the new form
$\mathcal{W}^{[\![2n]\!]}(t)=\sum_{\alpha\in\mathrm{Pair}[\![2n]\!]}\int_{s+s^{\prime}=t}\mathrm{d}s\,\,\mathcal{B}_{\alpha}(s)\,\mathcal{C}^{[\![2n]\!]}_{\alpha}(s^{\prime})\ .$ (3.19)
@@
[…]
@@
$\begin{split}\mathcal{W}^{[\![2n]\!]}(t):=\sum_{\vec{\alpha}\in\mathrm{Dgm}_{*}[\![2n]\!]}\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\,\mathcal{P}^{[\![2n]\!]}_{\alpha_{1}}(\tau_{1/2})^{*}\prod_{k=1}^{|\alpha|-1}&\mathcal{J}^{[\![2n]\!]}_{\alpha_{k}}(\tau_{k})\,\mathcal{P}^{[\![2n]\!]}_{\alpha_{k}\alpha_{k+1}}(\tau_{k+1/2})\\ &\cdot\mathcal{J}^{[\![2n]\!]}_{\alpha_{|\vec{\alpha}|}}(\tau_{|\vec{\alpha}|})\mathcal{P}^{[\![2n]\!]}_{\alpha_{|\vec{\alpha}|}}(\tau_{|\vec{\alpha}|+1/2}).\end{split}$ (3.22)
@@
Summing over $\vec{\eta}\in\mathrm{Dgm}_{\uparrow}[\![2n]\!]$ with $\vec{\eta}_{n}=\alpha$ being fixed gives $\mathcal{B}_{\alpha}$ in (3.18). Finally, read the rest of the integrand (after seeing the first operator that involves $\alpha_{k_{n}}=\eta_{n}=\alpha$), integrate over $\tau_{k}$ for $k\geq k_{n}$, and sum over $\alpha_{k}$ for $k>k_{n}$. We arrive at (3.19).
@@
$\|\mathcal{B}^{\ell}_{\vec{\alpha}}(\vec{s}\,)\|_{2\to q,a}\leq c\,(s_{1}\cdots s_{\ell})^{-1/p}\,(s_{2}\cdots s_{\ell})^{-1/2}\,e^{c\,(s_{1}+\ldots+s_{\ell})}\ .$ (3.31)
@@
The same bound holds in Case 2 by applying (2.10e) with $(\omega,\alpha)=(\eta_{\ell},\eta_{\ell})$ and (3.29) with $(\omega,\alpha)=(\eta_{\ell-1},\eta_{\ell-1})$ to (3.17). In (3.33), bounding $(s_{1}+\ldots+s_{\ell})^{-1/p}\leq s_{\ell}^{-1/p}$ and using the induction hypothesis conclude the bound (3.31). Next, use (3.31) for $\ell=n$ in (3.18). For the heat semigroup in (3.18), using (2.10a) followed by (2.10b) gives $|\langle f,\mathcal{P}^{\omega}(s)f^{\prime}\rangle|\leq c\,\|f\|_{p,a}\|f^{\prime}\|_{2}$, which implies $\|\mathcal{P}^{\omega}(s)\|_{2\to q,a}\leq c$. Hence
$\big{\|}\mathcal{B}_{\alpha}(t)\big{\|}_{2\to q,a}\leq c\,e^{ct}\sum_{\vec{\eta}\in\mathrm{Dgm}_{\uparrow}[\![2n]\!],|\eta_{n}|=\alpha}\int_{s_{1}+\ldots+s_{n}=t}\mathrm{d}s_{1}\cdots\mathrm{d}s_{n-1}(s_{1}\cdots s_{n})^{-1/p}\,(s_{2}\cdots s_{n})^{-1/2}\ .$ (3.34)
Evaluate this integral and bound the size of the sum by $c(n)=c$ (see (3.13)) to get $\|\mathcal{B}_{\alpha}(t)\|_{2\to q,a}\leq c\,e^{ct}t^{(n-1)/2-n/p}$. Insert this bound and (3.30) into (3.19), evaluate the integral, and bound the size of the sum by $c(n)=c$. Doing so gives the desired result (3.11).
4. Existence, proof of Proposition 1.4
We prove Proposition 1.4 in steps. In Section 4.1, we prepare the notation and tools. In Section 4.2, we check that every limit point of $\{Z^{\varepsilon}\}_{\varepsilon}$ satisfies Definition 1.2(b)(d). In Section 4.3, we prove that (the law of) $\{Z^{\varepsilon}\}_{\varepsilon}$ is tight in $C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},M_{+}(\mathbb{R}^{4}))$. Once these steps are completed, Proposition 1.4 follows from Theorem 1.3.

### v3

@@ 3.1. Lindeberg’s method
We use Lindeberg’s exchange method. Take $T=\{t_{1}<\ldots<t_{|T|}=1\}\subset G\cap(0,1]$, fix any $L^{2}$-applicable mollifier $\{u_{\ell}\}_{\ell}$, and let
$h_{i}(\ell,T)$ $:=r_{i-1}(\ell,T)-r_{i}(\ell,T)\ ,\qquad r_{i}(\ell,T):=\mathbf{E}\big[h\big(\widetilde{Z}_{0,t_{i}}\bullet_{u_{\ell}}Z_{t_{i},1}f\big)\big]\ ,$ (3.2)
with the convention that $t_{0}:=0$. We take $Z$ and $\widetilde{Z}$ to be independent within expectations like the one above. The product $\bullet_{u_{\ell}}$ above is well-defined thanks to Lemma 2.3. We then write
$\mathbf{E}\big[h(Z_{0,1}f)\big]-\mathbf{E}\big[h(\widetilde{Z}_{0,1}f)\big]=h_{1}(\ell,T)+\ldots+h_{|T|}(\ell,T)\ .$ (3.3)
Let $\operatorname{mesh}T:=\min\{t_{i}-t_{i-1}|i=1,\ldots,|T|\}$. We will send $\ell\to\infty$ first and $\operatorname{mesh}T\to 0$ second, where the second limit is possible thanks to the assumption that $G\subset\mathbb{R}$ is dense.
We seek to derive an approximation of $r_{i-1}(\ell,T)$ via the Taylor expansion. First, by Corollary 2.6(b) for $U^{\prime}=\{0,t_{i-1},t_{i},1\}$, we have $r_{i-1}(\ell,T)=\lim_{\ell^{\prime}\to\infty}\mathbf{E}[h\big(\widetilde{Z}_{0,t_{i-1}}\bullet_{u_{\ell}}Z_{t_{i-1},t_{i}}\bullet_{u_{\ell^{\prime}}}Z_{t_{i},1}f)]$. Further using Proposition 2.4 with $U^{\prime}=\{0,t_{i-1},t_{i},1\}$ and $V=\{t_{i-1},t_{i}\}$ gives
$\lim_{\ell\to\infty}r_{i-1}(\ell,T)=\lim_{(\ell,\ell^{\prime})\to(\infty,\infty)}\mathbf{E}\big[h\big(\widetilde{Z}_{0,s}\bullet_{u_{\ell}}Z_{s,t}\bullet_{u_{\ell^{\prime}}}Z_{t,1}f)\big]\Big|_{(s,t)=(t_{i-1},t_{i})}\ .$ (3.4)
@@
$\limsup_{\ell\to\infty}\Big|r_{i-1}(\ell,T)-\sum_{n=0}^{3}\frac{1}{n!}h^{n}_{t_{i-1},t_{i}}\Big|\leq c(h)\,v_{t_{i-1},t_{i}}\ .$ (3.8)
Next, we claim that (3.8) holds also for $r_{i}(\ell,T)$ in place of $r_{i-1}(\ell,T)$. Indeed, running the same argument above shows that (3.4)–(3.6) hold for $r_{i}(\ell,T)$ in place of $r_{i-1}(\ell,T)$, for $W_{s,t}$ in place of $\widetilde{W}_{s,t}$, and with $\ell$ and $\ell^{\prime}$ swapped. The swapping of $\ell$ and $\ell^{\prime}$ does not matter because the second limit in (3.4) is taken jointly in $\ell,\ell^{\prime}$. When evaluating the expectations in (3.6)–(3.7), we used only the moment formula (2.18) and the independence of $\widetilde{Z}_{0,s}$, $W_{s,t}$, $Z_{t,1}$. These properties hold also for $\widetilde{W}_{s,t}$ in place of $W_{s,t}$. Hence (3.8) holds for $r_{i}(\ell,T)$ in place of $r_{i-1}(\ell,T)$.
Let us bound (3.3). Combining (3.8) for $r_{i-1}(\ell,T)$ and for $r_{i}(\ell,T)$ gives
$\limsup_{\ell\to\infty}|h_{i}(\ell,T)|\leq c(h)\,v_{t_{i-1},t_{i}}\ .$ (3.9)
Next, in (3.7), use the symmetry of $\mathcal{Q}^{[\![4]\!]}(s)$ (its generator is self-adjoint) to write the inner product as $\langle\mathcal{Q}^{[\![4]\!]}(s)\bigotimes_{i=1}^{4}g_{i},\mathcal{W}^{[\![4]\!]}(t-s)\mathcal{Q}^{[\![4]\!]}(1-t)\bigotimes_{i=1}^{4}g^{\prime}_{i}\rangle$, and, for $p\in(2,\infty)$, use (2.16) to bound $\|\mathcal{Q}^{[\![4]\!]}(s)\bigotimes_{i=1}^{4}g_{i}\|_{p,1}$ and $\|\mathcal{Q}^{[\![4]\!]}(1-t)\bigotimes_{i=1}^{4}g^{\prime}_{i}\|_{p,1}$ by $c=c(p,\Gamma)$, where we encoded the dependence on $g_{i},g^{\prime}_{i}$ by $\Gamma$. The result gives $v_{s,t}\leq c(p,h,\Gamma)\|\mathcal{W}^{[\![4]\!]}(t-s)\|_{p,1\to q,1}$, where $1/p+1/q=1$. Inserting this bound into (3.9) and summing the result over $i$ give, for every $p\in(2,\infty)$,
$\big|\mathbf{E}\big[h(Z_{0,1}f)\big]-\mathbf{E}\big[h(\widetilde{Z}_{0,1}f)\big]\big|\leq c(p,h,\Gamma)\,\sum_{i=1}^{|T|}\big\|\mathcal{W}^{[\![4]\!]}(t_{i}-t_{i-1})\big\|_{p,1\to q,1}\ .$ (3.10)
The proof is completed upon applying Lemma 3.1 with $2n=4$ and any $p\in(6,\infty)$ and sending $\operatorname{mesh}T\to 0$. We prove the lemma in the next subsection.
@@
For any $2n\in 2\mathbb{Z}_{>0}$, $p\in(2,\infty)$, and $a,t\in(0,\infty)$, with $1/p+1/q=1$,
@@
The key to the proof is to derive a useful form of $\mathcal{W}^{[\![2n]\!]}(t)$. Fix $2n\in 2\mathbb{Z}_{>0}$, $p\in(2,\infty)$, and $a\in(0,\infty)$, write $c=c(n,p,a)$, and take $1/p+1/q=1$. Recall the expression of $\mathcal{W}^{[\![2n]\!]}(t)$ from (2.20) and also (2.12). The key to the proof is to rewrite that expression in a new form that exposes the nature of $\mathrm{Dgm}_{*}[\![2n]\!]$. To state the new form, we need some notation. For $\alpha\in\mathrm{Pair}(\omega)$, let
@@
$\mathcal{B}^{\ell}=\mathcal{B}^{\ell}_{\vec{\eta}}(\vec{s})=\mathcal{B}^{\ell}_{\eta_{1},\ldots,\eta_{\ell}}(s_{1},\ldots,s_{\ell}):L^{2}(\mathbb{R}^{2}\times\mathbb{R}^{2\omega_{\ell}\setminus\eta_{\ell}})\to L^{q,a}(\mathbb{R}^{2\omega_{\ell}})\$(3.14)
@@
Finally, define $\widetilde{\mathcal{B}}_{\vec{\eta}}(t):L^{2}(\mathbb{R}^{2[\![2n]\!]})\to L^{q,a}(\mathbb{R}^{2[\![2n]\!]})$ and $\widetilde{\mathcal{C}}_{\vec{\eta}}(t):L^{p,a}(\mathbb{R}^{2[\![2n]\!]})\to L^{2}(\mathbb{R}^{2[\![2n]\!]})$
$\widetilde{\mathcal{B}}_{\vec{\eta}}(t)$ $:=\int_{\Sigma(t)}\mathrm{d}\vec{s}\,\mathcal{B}^{n}_{\vec{\eta}}(s_{1},\ldots,s_{n})\otimes\mathcal{P}^{[\![2n]\!]\setminus(\eta_{1}\cup\cdots\cup\eta_{n})}(s_{1}+\ldots+s_{n})\ ,$ (3.18)
$\widetilde{\mathcal{C}}_{\vec{\eta}}(t)$ $:=\sum\int_{\Sigma(s)}\mathrm{d}\vec{\tau}\,\prod_{k=1}^{|\vec{\gamma}|-1}\mathcal{J}^{\omega}_{\gamma_{k}}(\tau_{k})\,\mathcal{P}^{\omega}_{\gamma_{k}\gamma_{k+1}}(\tau_{k+1/2})\,\cdot\mathcal{J}^{\omega}_{\gamma_{|\vec{\gamma}|}}(\tau_{|\vec{\gamma}|})\,\mathcal{P}^{\omega}_{\gamma_{|\vec{\gamma}|}}(\tau_{|\vec{\gamma}|+1/2})\ ,$ (3.19)
where the last sum runs over $\vec{\gamma}\in\mathrm{Dgm}[\![n]\!]$ under the constraint that $\eta_{1}\cup\cdots\eta_{n}\cup\gamma_{1}\cup\cdots\gamma_{|\gamma|}=[\![2n]\!]$. We will prove in the next few paragraphs that $\mathcal{W}^{[\![2n]\!]}(t)$ in (2.20) is equal to the new form
$\mathcal{W}^{[\![2n]\!]}(t)=\sum_{\vec{\eta}\in\mathrm{Dgm}_{\uparrow}[\![2n]\!]}\int_{s+s^{\prime}=t}\mathrm{d}s\,\,\widetilde{\mathcal{B}}_{\vec{\eta}}(s)\,\widetilde{\mathcal{C}}_{\vec{\eta}}(s^{\prime})\ .$ (3.20)
@@
For the second example, consider $\mathcal{P}^{123}_{12}(s_{1})^{*}\mathcal{J}^{123}_{12}(s_{2})\mathcal{P}^{123}_{12\ 23}(s_{3})$. Referring to (2.5a), we see that $\mathcal{P}^{123}_{12}(s_{1})^{*}=\mathcal{P}^{12}_{12}(s_{1})^{*}\otimes\mathsf{p}^{3}(s_{1})$, where $\mathsf{p}^{3}(s_{1})$ denotes the heat kernel acting on the 33rd coordinate. Hence, the diagram of $\mathcal{P}^{123}_{12}(s_{1})^{*}$ consists of the three leftmost single lines in Figure 3, with the top two merging. Similarly, the diagram $\mathcal{J}^{123}_{12}(s_{2})=\mathcal{J}^{12}_{12}(s_{1})\otimes\mathsf{p}^{3}(s_{1})$ consists of the double line and and the single line below it. As for $\mathcal{P}^{123}_{12\ 23}(s_{3})$, its diagram consists of the three rightmost single lines and has the effect of “swapping” the merged coordinates from $12$ to $23$. Note that the portion of Figure 3 within the dashed bubble represents
@@
$\begin{split}\mathcal{W}^{[\![2n]\!]}(t)=\sum_{\vec{\alpha}\in\mathrm{Dgm}_{*}[\![2n]\!]}\int_{\Sigma(t)}\mathrm{d}\vec{\tau}\,\mathcal{P}^{[\![2n]\!]}_{\alpha_{1}}(\tau_{1/2})^{*}\prod_{k=1}^{|\alpha|-1}&\mathcal{J}^{[\![2n]\!]}_{\alpha_{k}}(\tau_{k})\,\mathcal{P}^{[\![2n]\!]}_{\alpha_{k}\alpha_{k+1}}(\tau_{k+1/2})\\ &\cdot\mathcal{J}^{[\![2n]\!]}_{\alpha_{|\vec{\alpha}|}}(\tau_{|\vec{\alpha}|})\mathcal{P}^{[\![2n]\!]}_{\alpha_{|\vec{\alpha}|}}(\tau_{|\vec{\alpha}|+1/2}).\end{split}$ (3.23)
@@
Finally, read the rest of the integrand (after seeing the first operator that involves $\alpha_{k_{n}}=\eta_{n}=\alpha$), integrate over $\tau_{k}$ for $k\geq k_{n}$, and sum over $\alpha_{k}$ for $k>k_{n}$. We arrive at (3.20).
@@
$\|\mathcal{B}^{\ell}_{\vec{\eta}}(\vec{s}\,)\|_{2\to q,a}\leq c\,(s_{1}\cdots s_{\ell})^{-1/p}\,(s_{2}\cdots s_{\ell})^{-1/2}\,e^{c\,(s_{1}+\ldots+s_{\ell})}\ .$ (3.32)
@@
The same bound holds in Case 2 by applying (2.10e) with $(\omega,\alpha)=(\eta_{\ell},\eta_{\ell})$ and (3.30) with $(\omega,\alpha)=(\omega_{\ell-1},\eta_{\ell-1})$ to (3.17). In (3.34), bounding $(s_{1}+\ldots+s_{\ell})^{-1/p}\leq s_{\ell}^{-1/p}$ and using the induction hypothesis conclude the bound (3.32). Next, use (3.32) for $\ell=n$ in (3.18). For the heat semigroup in (3.18), using (2.10a) for $p=2$ followed by (2.10b) gives $|\langle f,\mathcal{P}^{\omega}(s)f^{\prime}\rangle|\leq c\,\|f\|_{p,a}\|f^{\prime}\|_{2}$, which implies $\|\mathcal{P}^{\omega}(s)\|_{2\to q,a}\leq c$. Hence,
$\big\|\widetilde{\mathcal{B}}_{\vec{\eta}}(t)\big\|_{2\to q,a}\leq c\,e^{ct}\int_{s_{1}+\ldots+s_{n}=t}\mathrm{d}s_{1}\cdots\mathrm{d}s_{n-1}(s_{1}\cdots s_{n})^{-1/p}\,(s_{2}\cdots s_{n})^{-1/2}\ .$ (3.35)
Evaluating this integral gives the bound $\leq c\,e^{ct}t^{(n-1)/2-n/p}$. Insert this bound and (3.31) into (3.20), evaluate the integral, and bound the size of the sum by $c(n)=c$ (see (3.13)). Doing so gives the desired result (3.11).
4. Convergence, proof of Proposition 1.5
We prove Proposition 1.5 in steps. In Section 4.1, we prepare the notation and tools. In Section 4.2, we check that every limit point of $\{Z^{\theta,\varepsilon}\}_{\varepsilon}$ satisfies Definition 1.2(b)(d). In Section 4.3, we prove that (the law of) $\{Z^{\theta,\varepsilon}\}_{\varepsilon}$ is tight in $C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},\mathcal{M}_{+}(\mathbb{R}^{4}))$. Once these steps are completed, Proposition 1.5 follows from Theorem 1.3.